Hi there
Hope you well and you can help me plse explain to me the solution tothis
2-3(N-6)+-4=(1/3+1/5)x1 7/8
Would really appreciate your help

Answer by checkley77(12569)

2-3(N-6)+-4=(1/3+1/5)x1 7/8
2-3N+18-4=[(5+3)/15]*15/8
-3N+16=(8/15)(15/8
-3N+16=1*1
-3N=-16+1
N=-15/-3
N=5 ANSWER.
PROOF:
2-3(5-6)-4=1
2-3*-1=1
2+3-4=1
5-4=1
1=1